OpenStudy (anonymous):

y varies directly as to x. when x = -1.5 y = 9. Find y when x = 6

OpenStudy (jdoe0001):

direct variaion y = \(\bf k\)x what's "k"? well, when x = -3/2, y = 9, thus 9 = \(\bf k\)(-3/2) solve for "k" to see what "k" is

OpenStudy (anonymous):

k is -3/2

OpenStudy (jdoe0001):

hmmm, let's see 9 = k(-3/2) what would that give you cross multiplying?

OpenStudy (anonymous):

-1.5

OpenStudy (jdoe0001):

\(\bf 9 =k \times -\cfrac{3}{2} \implies 9 = \cfrac{3k}{-2}\) try cross multiplying :)